OpenStudy (bookworm14):

@TheLukeskywalker2 @satellite73 Can you check my answer to this one please I will post it below--->

OpenStudy (bookworm14):

I think it has no solution but, im not sure I am working it out right... any help?

OpenStudy (anonymous):

lol of course it has a solution !

OpenStudy (anonymous):

just because you don't feel like finding it doesn't mean there isn't one...

OpenStudy (anonymous):

wanna find it?

OpenStudy (bookworm14):

I know but im supposed to determine if it has no solution 1 solution or infinite solutions and YES!!! I wanna find it !!!!!!!!!!!! :D

OpenStudy (anonymous):

ok first off since the coefficient of the x term in the left and right are different, there must be a solution one solution

OpenStudy (bookworm14):

this is the firsst ive done with the problem containing a fraction

OpenStudy (anonymous):

\[3-\frac{2}{3}x=-2x+5\] if you don't like the fractions, get rid of them you have an equal sign, multiply both sides by 3 and get \[9-2x=-6x+15\]

OpenStudy (anonymous):

notice you have to multiply EVERYTHING by 3

OpenStudy (bookworm14):

why do you multiply by 3?

OpenStudy (anonymous):

now we can solve without dealing with those nasty fractions

OpenStudy (bookworm14):

wait real quick can i solve it and you check my answer?

OpenStudy (anonymous):

take a look at \[3-\frac{2}{3}x=-2x+5\] and remember you don't like fractions now why would you multiply by 3?

OpenStudy (anonymous):

sure i will wait get a glass of juice (beer)

OpenStudy (bookworm14):

lol xD drink up

OpenStudy (bookworm14):

Ughhh nevermind I got stuck again \(\huge\color{blue}{\ddot\frown}\)

OpenStudy (anonymous):

\[9-2x=-6x+15\] add \(6x\) get \[9+4x=15\] subtract \(9\) get \[4x=6\] divide by \(4\) get \[x=\frac{6}{4}=\frac{3}{2}\]

OpenStudy (anonymous):

The only case where there are infinitely many solutions is when the equation reduces down to x=x, or something similar. The only way that there are no solutions is where you reduce to a mathematical contradiction, like 1=2. It is obvious that \[1\neq2\] so there are no values of x for which it would be true.

OpenStudy (anonymous):

\[\huge\color{blue}{\ddot\smile}\]

OpenStudy (anonymous):

\[\Huge\color{red}{\ddot\smile\clubsuit}\]

OpenStudy (anonymous):

Do you have any more Bookworm? If not I might be going to bed soon.

OpenStudy (bookworm14):

well both of those make sense \(\huge\color{purple}{Thank~You~!~!~!~! ~ \color{lime}{\Huge\ddot\smile \infty \heartsuit}}\) and Im okay for now, I think satelite73 will be here if i need anyhting else, sleep well :3 :) @TheLukeskywalker2

OpenStudy (anonymous):

I offer my assistance if needed as well Book

OpenStudy (bookworm14):

why thank you :3 @hilbertboy96

OpenStudy (anonymous):

no probs